# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ 3.3.2 Model-based RL
Both value-based and policy-based approaches in RL operate as model-free methods, meaning they do not utilize the model of the environment to plan ahead by simulating future steps. This is where MCTS Browne et al., (2012) comes into play, it is a heuristic search the combines the accuracy of tree search with the power of random sampling to efficiently explore large state spaces. The algorithm builds a search tree incrementally, where nodes represent states and edges represent actions.
The process begins with selection, where the algorithm chooses the most promising child nodes from the root until it reaches a leaf node. If the leaf node is not terminal, the expansion phase adds one or more child nodes. Next, during the simulation phase, the algorithm runs a simulation from these new nodes to a terminal state, typically using random actions. Finally, in backpropagation, the results of the simulation are used to update the values of the nodes in the path from the leaf to the root, propagating the success or failure of the simulation.
@@
AlphaZero, designed to master games like chess, shogi, and Go, utilizes deep neural networks combined with MCTS and relies on predefined game rules to achieve superhuman performance through self-play and RL Silver et al., (2016). MuZero extends this approach by generalizing to environments where the rules are not known in advance, integrating RL, MCTS, and learned models to predict environmental dynamics Schrittwieser et al., (2020). EfficientZero builds on MuZero and achieves super-human performance on the 100k Atari benchmark, surpassing previous state-of-the-art results significantly Ye et al., (2021). It introduces innovations like self-supervised consistency losses for accurate next-state predictions and end-to-end value prefix prediction to handle state aliasing issues. These enhancements improve exploration and action search capabilities, making EfficientZero highly effective in data-limited scenarios.
@@ 4 Graph Reinforcement Learning for Transmission Networks
Management and control are crucial issues in electricity transmission, requiring grid operators to meet all demand at all times. This task is becoming increasingly difficult due to rising demand and the variability from more renewable energy sources. Consequently, grid operation is shifting to adapt to generated power rather than predicted demand Viebahn et al., 2022b . Previously, grids were load-driven, adjusting generation based on estimated demand, but now traditional control methods are less effective as optimization problems grow more complex.
Traditional optimization methods face scalability issues with non-linear power systems, prompting a move towards more adaptive strategies. DRL is emerging as a possible solution for managing uncertainties and constraints in transmission grids Marot et al., (2021).
New grid operation approaches can’t be tested in real scenarios, so they are validated through simulations. Grid2Op Donnot, (2020) is the most commonly used environment, simulating grid operations and evaluating DRL agent performance across various scenarios. This environment is used in the Learning to Run a Power Network (L2RPN) Marot et al., (2020) challenges, which encourage the development of RL methods for grid monitoring and control.
Grid2Op Donnot, (2020) provides several actions, such as generator redispatch to balance supply and demand and curtailment of renewables to regulate output. A significant feature of the L2RPN challenges is real-time control of grid topology using bus switches and busbar splits Silver et al., (2017). This can avoid renewable energy curtailment and improve power transmission efficiency. Topological measures offer a low-cost and dynamic response to congestion and overload.
By training in realistic simulation environments, RL agents can enhance grid efficiency and resilience. However, the vast state and action spaces in transmission grids are challenging. Promising approaches to reduce action space involve heuristic methods Lehna et al., (2023); EI Innovation Lab, (2020); Chauhan et al., (2022) and hierarchical strategies that separate high-level topology decisions from specific actions Manczak et al., (2023); van der Sar et al., (2023); Yoon et al., (2021); Liu et al., (2024).
DRL solutions from these challenges have developed effective control strategies leveraging topological flexibility EI Innovation Lab, (2020); Lehna et al., (2023). Few approaches currently use GNNs, which are advantageous for topological actions by directly learning from the grid’s graph structure. Therefore, in this chapter we analyse existing GRL approaches. The success of DRL solutions in these competitions demonstrates their potential in power system operations. However, simulations can’t cover all aspects of real grid operations, and Grid2Op abstracts from reality by using standardized IEEE grids, potentially limiting real-world applicability. This focus may overlook unique issues of actual power grids.
We identified eight GRL approaches that address grid operation in transmission networks. Their primary objective is to maintain network stability under dynamic conditions in real-time. Thanks to a well-organized pre-dispatch schedule and the power system’s robustness, specific actions are typically unnecessary in routine situations. Consequently, all approaches activate agents only in critical states, such as heavily loaded lines. When no lines are heavily loaded, a "do-nothing" action is implemented, and the agent intervenes only when loading surpasses a threshold. This control mechanism prevents overacting and minimizes costs, as actions are associated with costs and should not occur when the grid is stable. This feature is common among most RL approaches EI Innovation Lab, (2020); Lehna et al., (2023) and GRL approaches Xu et al., (2020); Taha et al., (2022); Yoon et al., (2021); Sar et al., (2023). For example, Sar et al., (2023) employs a hierarchical RL strategy where the highest level agent is a rule-based mechanism triggered by critical line loading.
@@
Table 1: Overview of GRL approaches proposed for transmission grids. We categorize approaches by the used RL algorithm and GNNs, specifying action types and emphasizing topological actions alongside others. Grid sizes indicate the number of buses evaluated. The focus column describes distinctive aspects of each approach.
env Model Control Algorithm Action GNN Grid Size Focus
@@
Taha et al., (2022) MCTS for acion selection Topo GCN with residual connections IEEE 118 GCNs for power flow estimation
@@
Sar et al., (2023) Multi-Agent ISACD, DSACD & PPO Topo GCN IEEE 5 Hierarchical RL
@@ 4.1 RL Framework
Reward
When an action other than "do-nothing" is required, the agent is activated. The goal of the approaches analyzed here is to manage line flows, mitigate congestion, and minimize costs. The reward for GRL agents is based on these aspects. For example, Zhao et al., (2022) and Wu et al., 2023b combine operational costs like power loss and generator dispatch with penalties for constraint violations, such as voltage limits, battery state-of-charge (SoC) constraints, or line flow limits. Xu et al., (2020) use a reward considering line overflow proximity, a gameplay reward reflecting the successful completion of a scenario, and a reward for operational loss. Yoon et al., (2021), Qiu et al., (2022), and Sar et al., (2023) focus on grid efficiency, using the ratio of total served load to total production as a reward. This aims to reduce energy loss from congestion. In contrast, Taha et al., (2022) penalize line overloads with a binary reward counting time steps without overload. Similarly, Xu et al., 2022a aim to improve grid flexibility by penalizing heavily-loaded lines with appropriate weights, targeting potential transmission bottlenecks.
@@
In critical states, agents can choose to initiate redispatch or alter the grid’s topology by changing bus configurations or line connectivity. Often, lines are reconnected after being disconnected. Since bus-bar splitting by grid operators is cheaper than redispatch actions, it is preferred when possible. However, the action space for bus configurations is very large. These approaches are typically trained on the grid2op environment, which features double-busbar systems. This allows each of the $n$ buses to be set to one of two busbars, resulting in an exponential number of configurations per substation. Simulating each configuration is impractical, especially for larger grids. Even for the small IEEE 14-bus grid, 178 valid actions are available. Approaches like Xu et al., (2020) and Qiu et al., (2022) consider only do-nothing and topological actions. Grid2op restricts bus-bar splitting to one substation per time step, in line with L2RPN challenges and real grid operations. The specific actions considered by each analyzed approach are detailed in Tab 1.
@@
Most approaches use the information available in the grid2op environment, though not all use every available feature. For example, Xu et al., (2020) include grid topology, connected elements, and features of generators, loads, and power lines. This data includes the bus-bar index, active and reactive production and consumption, voltage set points and magnitudes, and the active and reactive flow for each power line. Additionally, they use the ratio ($\rho$) between current flow and thermal limits. While state information is generally consistent across GRL approaches, grid size varies by study, affecting the state size. Only Wu et al., 2023b , operating in a different environment, use just voltages in their state representation. Typically, states are modeled as graphs and embedded using GNN approaches, which are described in subsection 4.3 along with the corresponding GNNs.
@@ 4.2 Overall Approach and RL Algorithms
While the states, actions, and rewards of the analyzed GRL frameworks are similar, they use different RL algorithms and focus on various problem aspects. Xu et al., (2020), using Double-Q-learning, note that conventional RL agents often violate constraints during exploration. To mitigate this, they employ a soft constraint, replacing actions that violate basic constraints with "do-nothing". During exploitation, they employ an adaptive strategy to select the top $N$ actions with the highest Q-values, checked by the Grid2op simulation to avoid catastrophic results. This strategy improves action selection during early training and saves time when the agent is well-trained.
@@
Instead of Grid2op simulations, Taha et al., (2022) use a GNN to predict the lines loading $\rho$ for different topologies. They estimate state evolution under various actions and select the trajectory maximizing cumulative rewards using MCTS. Starting with the root as the initial state and edges representing actions leading to subsequent nodes, a tree is built iteratively, keeping only nodes with low loads. The above-mentioned GNN predicts the line loading under the corresponding action. After simulating steps with a do-nothing agent, they select the optimal actions by maximizing node values and potential future actions.
@@
Another MCTS-based approach is proposed by Xu et al., 2022a . Their tree structure resembles that of Taha et al., (2022), with leaf nodes representing subsequent overload states from feasible actions. Like Taha et al., (2022), they determine the best action based on the path yielding the highest value. They employ double-dueling Q-networks, training multiple parallel sub-agents to select actions from distinct sub-action spaces derived from MCTS. They use a long-short term strategy to balance immediate and long-term benefits and manage sub-agents effectively. Each agent simulates $n$ actions at overload states, selecting the best action through efficient comparison. They also introduce a constraint to limit topological actions and keep redispatch within physically feasible limits.
@@
In contrast to other approaches, Yoon et al., (2021) focus less on extensive simulation and instead develop an afterstate representation that captures the deterministic grid topology after applying a topological action. This approach directs their RL algorithm towards understanding the stochastic dynamics following each action. They utilize a hierarchical policy with a high-level policy generating desired topologies and a low-level policy executing topology changes. This strategy avoids the need to learn individual actions by focusing on appropriate topologies for critical situations. They explore rule-based approaches, such as CAPA, to efficiently sequence actions, prioritizing substations with high-capacity usage and ensuring timely responses. Their actor-critic algorithm uses the afterstate representation and goal topology predictions to enhance exploration and improve value function determination. Similarly, Qiu et al., (2022) adopt an approach resembling Yoon et al., (2021), employing a different attention mechanism in their Graph Convolutional Network (GCN).
In a novel approach, Sar et al., (2023) employ a 3-level agent system to manage complex decision-making in transmission grids. The highest-level agent assesses the need for action at each time step, often opting for inaction in safe conditions. It uses a rule-based method to identify critically loaded lines. In unsafe scenarios, this agent activates the mid-level agent, which prioritizes substations needing action based on CAPA Yoon et al., (2021), focusing on substations with high line loads. At the lowest level, substation-specific agents manage individual substations by selecting bus assignments within a predefined action space, excluding symmetric actions. This approach stands out for its use of multiple substation-specific agents and the comparison of parameter-shared critics versus independent agents in PPO and SAC.
Unlike other approaches focusing on topological changes, Zhao et al., (2022) tackle the variability in power system topologies caused by factors like extreme weather events and maintenance plans. They emphasize stability and resilience through re-dispatching, curtailment, and battery storage actions rather than altering the grid’s structure. Similar to EI Innovation Lab, (2020), they employ imitation learning to pre-train a PPO RL agent, using a GNN based on GraphSAGE Hamilton et al., 2017b to handle dynamic power network topologies. This approach demonstrates adaptability to topology changes due to emergencies, maintenance, and new bus constructions.
@@
The approach by Wu et al., 2023b tackles a broad problem involving stochastic dynamic OPF with Renewable Energy Sources (RES) and decentralized energy systems, which present significant computational challenges. Their method employs an actor-critic framework where the actor, updated based on the critic’s assessment, anticipates future actions to ensure viability of current decisions. Separate neural networks predict voltages, and critic networks are refined using temporal difference learning. The framework integrates problem constraints using Lagrangian multipliers, leveraging the duality principle to optimize both primal and dual variables through gradient-based primal dual updates. This concept is also applied in similar constrained RL problems like Yan et al., (2023).
@@ 4.3 Graph Embeddings
Looking at the representations of the states in these GRL approaches, the grid state is represented as a graph to encompass its elements. Utilizing GNNs, these methods extract features to learn a latent representation of the grid incorporating both its features and topology. While the overall procedure remains consistent across approaches, variations exist in how the power grid is modeled and in the specific GNN architectures used for feature extraction.
Grid2opDonnot, (2020) provides a graph representation where each load, generator, and endpoint of a transmission line is a node. Nodes are connected based on busbar assignments within substations. Each transmission line is represented by two (connected) nodes to account for potential connections to different busbars. This approach simplifies scenarios where substations are split, ensuring consistent handling across different configurations. Node features vary slightly across approaches, typically including all available variables from grid2op, with missing features padded with zeros as necessary.
@@
Xu et al., (2020) utilize the grid2op graph representation and employ a spectral GNN that uses spectral graph convolution to process information within this grid topology. This spectral GNN effectively extracts features from irregular non-Euclidean data. On the other hand, Taha et al., (2022) employ two adjacency matrices to indicate connections at both ends of each line. Each power line is associated with a feature vector that aggregates injections from nodes connected to the same bus at each end. Their processing GNN predicts system dynamics using a structure comprising 2 fully connected layers and 3 graph convolutional layers with residual connections. Following this, two additional fully connected layers make the final predictions. Notably, Taha et al., (2022) use GNNs to predict relative line loads ($\rho$) for an ad-hoc MCTS method. They incorporate a critical step where the GNN is trained on topologies resembling a reference topology. When the GNN’s generalization error increases, they revert the grid’s topology back to this reference state, which has shown positive effects in maintaining grid stability, as affirmed by Lehna et al., (2023, 2024).
@@
Conversely, approaches like Xu et al., (2020); Yoon et al., (2021) employ GNNs directly for learning policy and value functions in DRL contexts. For instance, Yoon et al., (2021) utilize a GNN that comprises transformer-based attention layers for node embeddings, shared across actor and critic heads. Actor heads output mean and standard deviation parameters for action sampling from a normal distribution. Critics similarly process node embeddings to derive scalar values representing value functions. They focus on controllable substations, typically those with more than two elements, to simplify the goal space dimensionality.
@@
In another study on smaller grids like IEEE 5, Sar et al., (2023) utilize three GNN blocks for shared layers in both single and multi-agent settings. The actor incorporates three GNN blocks, while the critic employs one. Although the exact input graph structure isn’t detailed, Grid2op’s typical graph model is assumed.
@@
Meanwhile, Zhao et al., (2022) address dynamic topologies using GraphSAGE within PPO networks to adapt to evolving grid characteristics. GraphSAGE’s unsupervised learning generates node embeddings independently from PPO losses, facilitating generalization to unseen nodes via aggregator functions. It leverages neighborhood sampling, such as random walks, to generate embeddings in changing topologies with varying connectivities and node counts. Their PPO agent integrates GraphSAGE-trained networks with fully connected layers, with shared components between actor and critic, differing only in output for actions and state values.
@@
On a different note, Xu et al., 2022a integrate GATs including multiple attention heads to process power grid states, enabling better adaptation to unpredictable topological changes and improve generalization across different structures. The GAT layers serve as models for both the actor and critic component of PPO, followed by Dense layers to output discrete actions for each agent.
@@
Wu et al., 2023b utilize a Chebyshev Spatio-Temporal Graph Convolutional Network (STGCN) to model dynamics over graph-structured data, specifically tailored for analyzing power grid topologies. The architecture combines temporal convolutional layers with spatial convolutional layers to capture temporal dependencies and spatial interactions within the grid. Temporal convolutional layers employ 1-D CNNs with specified kernel widths and output channels to capture temporal evolution, while spatial convolutional layers extract spatial features using complex-valued transfer functions. This integrated approach yields multiple output channels at each layer.

### v3

@@ 3.3.2 Model-based RL
Both value-based and policy-based approaches in RL operate as model-free methods, meaning they do not utilize the model of the environment to plan ahead by simulating future steps. This is where MCTS Browne et al., (2012) comes into play; it is a heuristic search that combines the accuracy of tree search with the power of random sampling to efficiently explore large state spaces. The algorithm builds a search tree incrementally, where nodes represent states and edges represent actions.
The process begins with selection, where the algorithm chooses the most promising child nodes from the root until it reaches a leaf node. If the leaf node is not terminal, the expansion phase adds one or more child nodes. Next, the algorithm runs a simulation from these new nodes to a terminal state, typically using random actions. Finally, in backpropagation, the results of the simulation are used to update the values of the nodes in the path from the leaf to the root, propagating the success or failure of the simulation.
@@
Designed to master games such as chess, shogi and go, AlphaZero uses deep neural networks combined with MCTS and relies on predefined game rules. It learns by playing and RL Silver et al., (2016). MuZero extends this approach by generalizing to environments with unknown rules, integrating RL, MCTS, and learned models to predict environmental dynamics Schrittwieser et al., (2020). EfficientZero builds on MuZero and achieves superhuman performance on the 100k Atari benchmark, significantly outperforming previous state-of-the-art results Ye et al., (2021). It introduces innovations such as self-supervised consistency losses for accurate next-state prediction and end-to-end value prefix prediction to deal with state aliasing issues. These enhancements improve exploration and action search capabilities, making EfficientZero highly effective in data-limited scenarios.
@@ 4 Graph Reinforcement Learning for Transmission Networks
To ensure safe and reliable transmission, human experts manually manage power grids. However, the rise in renewable energy and demand necessitates automated, data-driven optimization Marot et al., (2021), shifting grid operation to adapt to generated power rather than predicted demand Viebahn et al., 2022b .
Many transmission grid control methods focus on generation or loads, like re-dispatch Kamel et al., (2020); Bai et al., (2023); Fuxjäger et al., (2023) or load shedding Larik et al., (2019). Topology actions, such as bus switches and busbar splits Silver et al., (2017), offer a cost-effective alternative, enabling efficient power flow rerouting using the existing infrastructure. Despite the non-linear and large-scale nature of the problem, there are some promising approaches to reduce action spaces that involve heuristic methods Lehna et al., (2023); EI Innovation Lab, (2020); Chauhan et al., (2022) and hierarchical strategies Manczak et al., (2023); van der Sar et al., (2023); Yoon et al., (2021); Liu et al., (2024).
New grid operation approaches are validated through simulations, often using the Grid2op Donnot, (2020) environment from the Learning to Run a Power Network (L2RPN) challenges Marot et al., (2020). Grid2Op is crucial for developing methods to tackle congestion and enhance grid reliability by simulating operations and evaluating DRL agents’ performance across various scenarios. However, simulations cannot cover all aspects of real grid operations, and Grid2Op abstracts from reality by using standardized IEEE grids, potentially limiting real-world applicability. This focus may overlook unique issues of actual power grids.
We identified eight GRL approaches for managing transmission grids and maintaining stability under dynamic conditions in real-time. Due to a robust pre-dispatch schedule, actions are necessary only in critical states. In stable conditions, agents do not act ("do-nothing" action) and intervene only when the line loading exceeds a threshold. This minimizes costs by preventing unnecessary actions. This procedure is common among RL and GRL approaches EI Innovation Lab, (2020); Lehna et al., (2023); Lehna et al., 2024b ; Taha et al., (2022); Yoon et al., (2021); Sar et al., (2023). Tab 1 lists the RL method, action type, GNN architecture, grid size, and overall focus of the analyzed approaches.
@@
Table 1: Overview of GRL approaches proposed for transmission grids We categorize approaches by the used RL algorithm and GNNs, specifying action types and emphasizing topological actions (Topo) alongside others. Grid sizes indicate the number of buses evaluated. The focus column describes distinctive aspects of each approach.
env Approach Control Algorithm Action GNN Grid Size Focus
@@
Taha et al., (2022) MCTS for action selection Topo GCN with residual connections IEEE 118 GCNs for power flow estimation
@@
Sar et al., (2023) Multi-Agent SACD& PPO Topo GCN IEEE 5 Hierarchical RL
@@ 4.1 RL Framework
Rewards
The rewards are based on managing line flows, mitigating congestion, and minimizing costs. Examples include combining operational costs like power loss and generator dispatch with penalties for constraint violations such as voltage or line flow limits Zhao et al., (2022); Wu et al., 2023b and focusing on grid efficiency by rewarding the ratio of generated to served electricity to reduce energy loss due to congestion Yoon et al., (2021); Qiu et al., (2022); Sar et al., (2023). Similarly, Xu et al., (2020) use rewards for line overflow proximity, scenario completion, and operational loss. Others target transmission bottlenecks by penalizing heavily loaded lines Xu et al., 2022a ; Taha et al., (2022).
@@
In critical states, agents can re-dispatch or alter the grid topology by changing bus configurations or line connectivity, often reconnecting disconnected lines. The actions considered by each approach are listed in Tab. 1. For economic reasons, topological actions are preferred. Grid2op’s double-busbar system results in a large action space, making it impractical to simulate every configuration in larger grids.
@@
Most approaches use Grid2op information, although not all features are used. The state typically includes grid topology, connected elements, and features such as bus-bar data, generation and loads, voltages, and line flows. Furthermore, the ratio ($\rho$) between the current flow and the thermal limit of each line is a critical feature. Typically, the states are modeled as graphs embedded using GNNs (see 4.3). The state information is mostly consistent across the GRL approaches, only Wu et al., 2023b operate in a different environment and use only voltages as states.
@@ 4.2 Overall Approach and RL Algorithms
While the states, actions, and rewards of the analyzed GRL frameworks are similar, they utilize different RL algorithms. Xu et al., (2020), using Double-Q-learning, note that conventional RL agents often violate constraints during exploration. To mitigate this, they use a soft constraint to replace violating actions with "do-nothing". During exploitation, an adaptive strategy selects the top $N$ actions with the highest Q-values and verifies them using the Grid2op simulation to avoid catastrophic results. This improves action selection early in training and saves time when the agent is well-trained.
@@
Instead of Grid2op simulations, Taha et al., (2022) use a GNN to predict the line loading $\rho$ for different topologies. They estimate the state evolution under various actions and select the trajectory that maximizes the cumulative rewards using MCTS. Starting with the initial state as the root and actions as edges that lead to subsequent nodes, they iteratively build a tree, retaining nodes with low loads. The GNN predicts line loading for each action. After simulating steps with a do-nothing agent, they select optimal actions by maximizing node values and the number of possible future actions.
@@
Xu et al., 2022a present another MCTS-based approach with a similar tree structure like Taha et al., (2022). The leaf nodes represent overload states from feasible actions and the best action is determined by the highest-value path. Using double-dueling Q-networks, they train multiple sub-agents to select actions from different sub-action spaces derived from MCTS. A long-short-term strategy balances immediate and long-term benefits, managing sub-agents effectively. Each agent simulates $n$ actions at overload states, selecting the best through efficient comparison. They also constrain topological actions and re-dispatch to stay within feasible limits.
@@
Rather than using simulations like the above approaches, Yoon et al., (2021) develop an afterstate representation to capture grid topology after a topological action. This directs their RL algorithm to understand the stochastic dynamics following each action. They use a hierarchical policy: a high-level policy generates desired topologies, and a low-level policy executes changes. This strategy avoids learning individual actions by focusing on topologies for critical situations. They use rule-based approaches like CAPA to prioritize substations with high-capacity usage and ensure timely responses. Their actor-critic algorithm enhances exploration and value function determination using the afterstate representation and goal topology predictions. Qiu et al., (2022) use a very similar approach with a different attention mechanism in their Graph Convolutional Network (GCN).
Sar et al., (2023) present a novel three-level approach where the top-level agent decides on the need for action and identifies critically loaded lines. It activates the mid-level agent in unsafe scenarios and prioritizes high-load substations using CAPA Yoon et al., (2021). At the lowest level, substation-specific agents select bus assignments from a predefined action space. This approach is noteworthy for comparing parameter-shared versus independent critics in PPO and SAC.
To address the variability in grid topologies caused by, e.g., extreme weather or maintenance, Zhao et al., (2022) focus on re-dispatching, curtailment, and battery storage to ensure stability. Similar to EI Innovation Lab, (2020), they use imitation learning to pre-train a PPO agent with a GNN based on GraphSAGE Hamilton et al., 2017b .
@@
Wu et al., 2023b address the computationally challenging problem of stochastic dynamic OPF with Renewable Energy Sources (RES) and decentralized energy systems. They use an actor-critic method where separate neural networks predict voltages. The critic networks are refined with temporal difference learning. They integrate constraints with Lagrangian multipliers, leveraging the duality principle to optimize both primal and dual variables through gradient-based updates. This approach is also used in similar constrained RL problems, such as in Yan et al., (2023).
@@ 4.3 Graph Embeddings
In all analyzed approaches, the state is modeled as a graph from which a GNN extracts features. The resulting latent representation of the grid then contains both its features and its topology. All approaches follow this procedure, but they differ in the way the grid is modeled and in the specific GNN architectures used for feature extraction.
The Grid2op Donnot, (2020) environment models loads, generators, and transmission lines as nodes that are connected according to busbar assignments within substations. Each transmission line endpoint is represented by one node. This accounts for potential connections to different busbars and simplifies handling substations that are split. Node features generally include the available state information. Xu et al., (2020) utilize exactly this graph representation and embed it using a basic spectral GNN as defined in 6.
@@
In contrast, Taha et al., (2022) use two adjacency matrices to represent connections at both ends of a line. The node features vector aggregates injections from nodes connected to the same bus at each line end. Their GNN predicts system dynamics using three graph convolutional layers with residual connections combined with 2 fully connected neural network blocks. It predicts line loads ($\rho$) for an ad-hoc MCTS method and is trained on topologies similar to a reference topology. If the GNN shows increased generalization error, they revert to the reference topology, which helps maintain grid stability, as supported by Lehna et al., (2023); Lehna et al., 2024b .
@@
Conversely, Xu et al., (2020) and Yoon et al., (2021) use GNNs for learning policy and value functions in RL. Yoon et al., (2021) employ a GNN with transformer-based attention layers for node embeddings, shared across actor and critic heads. The actor outputs parameters for action sampling from a normal distribution, while the critic derives value functions from node embeddings. They consider only substations with more than two elements to reduce the action space.
@@
In another study on smaller grids such as IEEE 5, Sar et al., (2023) apply three GNN blocks shared in both single and multi-agent settings. The actor consists of three GNN blocks, while the critic employs only one. Since the exact graph representation isn’t detailed, the Grid2Op graph described above is assumed.
@@
Meanwhile, Zhao et al., (2022) use GraphSAGE within PPO networks. This GNN generates node embeddings independent of PPO losses, improving generalization to topologies through neighborhood sampling. Their PPO agent combines GraphSAGE-trained networks with fully connected layers that share components between actor and critic but differ in outputs for actions and state values.
@@
On a different note, Xu et al., 2022a use GATs with multiple attention heads to handle power grid states, improving adaptation to unpredictable topological changes and generalization across structures. The GAT layers are used for both the actor and critic in PPO, with Dense layers generating discrete actions for each agent.
@@
Wu et al., 2023b use a Chebyshev Spatio-Temporal Graph Convolutional Network (STGCN) to model dynamics in power grid topologies. Their architecture combines temporal convolutional layers with spatial convolutional layers to capture both temporal and spatial dependencies. Temporal layers use 1-D CNNs with specified kernel widths and output channels, while spatial layers use complex-valued transfer functions to extract features, resulting in multiple output channels at each layer.